You throw a ball vertically at a height H. What height does the ball have to reach for it to spend twice as long in the air? (Note: you need to get an equation)
Answer #1

Suppose the ball was thrown at a velocity Vi then where v is final velocity and h is height.
  g is the acceleration due to gravity

Now at the maximum height h -H and vf 0

When the ball will come down from height H uf = 0-2gH initial velocity (vi) will be 0

I f the time taken to reach height H is ti then, rom Uf = Vi-gt] 2H


If the time taken to fall down from height H is t2 then [from uf = Vi + gt 2H to

So the ball takes same time to reach a height H and to fall down from the same height
Therefore each time will be twice to spend twice as long in the air If new initial velocity be vi then, 2H ひ

Now suppo se will reach a height h spending twice as long in the air h24H


The ball have to reach a height 4H to spend twice as long in the air